hello find the slo, x & y inercept.
please show me how so I can do the next 10 problems. Thanks
y=5

Answer by queenofit(120)

to find the slope you need some points, you didn't give any so I will give an example 1). (2,3) and 2). (2,4)
you actually can get the points by making a T chart. let x equal at least 5 different points and put those values you chose into your equation for x and solve for y. or do it backwards and solve for x
when you have points ( my example of points above) you put the points into the following equation and you get your slope
y2-y1
------= slope the 1 and 2 are just used to show you which set of points
x2-x1
so for my example of points, to find the slope you would
4-3
----=1 one would be your slope
2-2
now you need to put that in point slope form
y-y1=m(x-x1)
y-3= 1(x-2)
y-3=x-2
y=x+3-2
y=1x+1 this is y=mx+b form and the equatio of the line
use this to get your y and x intercept
set x=0 to get the y intercept and y=o to get the x intercept
0=x+1
-1=x x intercept would be (-1,0)
y=0+1
y=1 y intercept would be (0,1)
does this help any? I didn't know what you meant really, but this is how you find the slope and the y and x intercept.